# Data-Driven Superstabilizing Control under Quadratically-Bounded Errors-in-Variables Noise

### v1

@@ VI Numerical Example
$A_{\star}=\begin{bmatrix}0.6863&0.3968\\ 0.3456&1.0388\end{bmatrix},\ B_{\star}=\begin{bmatrix}0.4170&0.0001\\ 0.7203&0.3023\end{bmatrix}.$ (54)
Data $\mathcal{D}$ is collected from an execution of (54) over a time-horizon of $T=14$. In the data collection, the EIV noise $\Delta x,\Delta u$ are i.i.d. distributed according to $\Delta x_{t}\in\mathcal{N}(\mathbf{0},0.04^{2}I_{2}),\Delta u_{t}\in\mathcal{N}(\mathbf{0},0.03^{2}I_{2})$. It is desired to create a superstabilizing $(W=I_{2})$ controller that will succeed in regulating the ground-truth system with joint probability $P{{}_{\textrm{joint}}}=95\%.$ The per-noise probability is chosen as $\delta_{x}=\delta_{u}=(0.95)^{1/(2T-1)}=0.9981$. The $(\delta_{x},\delta_{u})$ chance constraint is modeled as $\lVert\Delta x_{t}\rVert_{2}\leq 0.04\epsilon(\delta_{x};2)=0.1416$ and $\lVert\Delta u_{t}\rVert_{2}\leq 0.03\epsilon(\delta_{u};2)=0.1062$ for each $t$.
Superstabilizing control is performed under a decay-bound objective $\lambda^{*}=\min_{\lambda,K}\lambda\geq\lVert A+BK\rVert_{\infty}\ \forall(A,B)\in\mathcal{P}$. The $k=1$ truncation of the Alternatives certification (IV-A2) (dense, $p_{A}=27$) returns the following policy:
$K_{\textrm{dense}}$ $=\begin{bmatrix}-1.1559&-0.1193\\ 1.1333&-0.9109\end{bmatrix}$ $\lambda^{*}_{\textrm{dense}}=0.9907.$ (57)
Worst-case superstabilization is certified by the value $\lambda^{*}<1$. The ground-truth system (54) has a decay-bound of $\lambda_{\star}=\lVert A_{\star}+B_{\star}K_{\textrm{dense}}\rVert_{\infty}=0.8219$ with respect to the controller $K$ in (57). The Full program ($p_{F}=62$) and Sparse ($p_{B}=18$) fail to perform superstabilization at $k=1$ because the returned bound satisfies $\lambda^{*}_{\textrm{full}}=3.1749\times 10^{3},\lambda^{*}_{\textrm{sparse}}=2.4173\geq 1$. Attempting execution of the $k=2$ tightening for the Full program ($p_{F}=\binom{64}{2}=2016$) results in out-of-memory errors in Mosek. For this specific example, the algorithms from Theorems 1 and 2 of [19] both fail to find a common quadratically stabilizing controller on (matrix-ellipsoid-based over-approximations of) $\mathcal{P}$.
@@ VII Conclusion
This paper presented a solution approach for data-driven superstabilization in the quadratically-bounded EIV setting. The $W$-superstabilization problem was formulated as an infinite-dimensional linear program, and was discretized using SOS-matrix methods. A Theorem of Alternative was used to eliminate the $(\Delta x,\Delta u)$ noise terms, resulting in matrix SOS constraints with lower computational complexity. Future work involves decreasing the runtime involved in posing solutions to EIV superstabilization, such as by investigating other approaches for discretization of the infinite-dimensional LPs. Other aspects include incorporating streaming data for online superstabilization, which could include integrating the EIV method into model predictive control.

### v2

@@ VI-A White Noise
$A_{\star}=\begin{bmatrix}0.6863&0.3968\\ -0.3456&1.0388\end{bmatrix},\ B_{\star}=\begin{bmatrix}0.4170&0\\ 0.7203&-0.3023\end{bmatrix}.$ (32)
Data $\mathcal{D}$ is collected from an execution of (32) over a time-horizon of $T=13$. In the data collection, the EIV noise $\Delta x,\Delta u$ are i.i.d. distributed according to $\Delta x_{t}\sim\mathcal{N}(\mathbf{0},0.03^{2}I_{2}),\Delta u_{t}\sim\mathcal{N}(\mathbf{0},0.025^{2}I_{2})$. It is desired to create an extended superstabilizing controller that will succeed in regulating the ground-truth system with joint probability $P{{}_{\textrm{joint}}}=95\%.$ The per-noise probability is chosen as $\delta_{x}=\delta_{u}=(0.95)^{1/(2T-1)}=0.9981$. The $(\delta_{x},\delta_{u})$ chance constraint is modeled as $\lVert\Delta x_{t}\rVert_{2}\leq 0.03\epsilon(\delta_{x};2)=0.1056$ and $\lVert\Delta u_{t}\rVert_{2}\leq 0.025\epsilon(\delta_{u};2)=0.08796$ for each $t$.
Extended superstabilizing control is performed to minimize $\lambda$ such that $\forall i:\lambda\geq v_{i}$. The $k=1$ truncation of (31) using the dense Alternatives method (26) ($p_{A}=27$) returns
$K_{\textrm{dense}}$ $=\begin{bmatrix}-0.9000&-0.8807\\ 0.1564&0.3679\end{bmatrix}$ $v=\begin{bmatrix}0.5519\\ 1.4481\end{bmatrix}.$ (33)
The Full (15), Dense (26), and Sparse (28) superstabilization $(W=I)$ programs are all infeasible for $k=1$. Attempting execution of the $k=2$ tightening for the Full program ($p_{F}=\binom{64}{2}=2016$) results in out-of-memory errors in Mosek. For this specific example, the algorithms from Theorems 1 and 2 of [19] both fail to find a common quadratically stabilizing controller.
VI-B Monte Carlo Test
This second example involves a Monte Carlo test for robust stabilization of 300 randomly generated 2-state 2-input ground-truth plants. Elements of the plant matrices $(A,B)$ are each i.i.d. drawn from a unit normal distribution. The noise $\Delta x_{t}$ and $\Delta u_{t}$ are drawn i.i.d. uniformly from unit $L_{2}$-balls of radius $\epsilon_{x}=0.225$ and $\epsilon_{u}=0.1$ respectively for a time horizon of $T=14$. At $k=1$, dense and sparse superstable SOS restrictions stabilized 48 and 42 systems respectively. Dense and sparse extended superstabilization $k=1$ SOS restrictions stabilized 71 and 62 systems respectively. Theorem 1 of [19] was infeasible at each instance. Theorem 2 of [19] stabilized 61 systems, with an overlap of 31 stabilized systems with the $k=1$ dense superstability, and 38 systems with the $k=1$ dense extended superstability.
@@ VII Conclusion
This paper presented a solution approach for data-driven superstabilization in the quadratically-bounded EIV setting. The $W$-superstabilization problem was formulated as an infinite-dimensional linear program, and was discretized using SOS-matrix methods. A Theorem of Alternative was used to eliminate the $(\Delta x,\Delta u)$ noise terms, resulting in matrix SOS constraints with lower computational complexity.
Future work involves reducing conservatism of EIV-aware control methods and incorporating streaming data for EIV-tolerant model predictive control.